OpenStudy (anonymous):

(i)Show that cot 75 degree = 2- sqrt 3 (ii) Hence, or otherwise show that cosec^2 75 degree = 4 cot 75 degree

OpenStudy (anonymous):

firstly, cot=cos/sin. and after all, you can use the trigo formula. split 75=45+30 so cos(30+45)=cos45cos30-sin45sin30 sin(30+45)=sin30cos45+sin45cos30

OpenStudy (anonymous):

Thanks. how to solve for part ii?

OpenStudy (anonymous):

cos(30+45)=cos45cos30-sin45sin30=sqrt2/2*sqrt3/2-sqrt2/2*1/2 sin(30+45)=sin30cos45+sin45cos30=1/2*sqrt2/2+sqrt2/2*sqrt3/2

OpenStudy (anonymous):

so you need to use, cos(30+45)/sin(30+45)=(sqrt2+sqrt6)/(sqrt6-sqrt2) (simplified, both numerator and denominator times 4) so now, you can times (sqrt2 + sqrt6) for both numerator and denominator, it will give you (sqrt2 + sqrt6)^2/4 which is 2+ sqrt3

OpenStudy (anonymous):

i thought is minus?

OpenStudy (anonymous):

oh, yest. just a typo. so you need to use, cos(30+45)/sin(30+45)=(sqrt6-sqrt2)/(sqrt6+sqrt2) so now, you can times (sqrt2 - sqrt6) for both numerator and denominator, it will give you (sqrt2 - sqrt6)^2/4 which is 2+ sqrt3

OpenStudy (anonymous):

which gives you 2- sqrt3

OpenStudy (anonymous):

for the second part, we all know that cosec^2=1 + cot^2. So just sub in the formula, 4cot75= 4(2-sqrt3)=8-4sqrt3 1+cot^2=1 + (2-sqrt3)^2=1+4+3-4sqrt3=8-4sqrt3

OpenStudy (anonymous):

for the second part, we all know that cosec^2=1 + cot^2. So just sub in the formula, 4cot75= 4(2-sqrt3)=8-4sqrt3 1+cot75^2=1 + (2-sqrt3)^2=1+4+3-4sqrt3=8-4sqrt3

OpenStudy (anonymous):

Thanks fireofsky.